# A Neutron Sensitive Detector Using 3D-Printed Scintillators

### v1

@@ Abstract
This work reports on the performance of a novel neutron-sensitive scintillating detector fabricated using Fused-Deposition Modelling (FDM) additive manufacturing. FDM is a cost-effective 3D-printing method employing flexible plastic filaments to create custom-shaped components. Scintillating filaments, based on polystyrene doped with p-terphenyl and 1,4-bis (5-phenyloxazol-2-yl) benzene, and enriched with 6LiF to enable neutron sensitivity were manufactured in house and achieved visible scintillation with a light output of 305 photons per MeV. Printed scintillators were then integrated into a detector system consisting of an image intensified TimePix3 camera, offering high spatial and temporal resolution. The detector performance was compared with Geant4 simulations of the scintillating sensor’s response to electrons, gamma-rays, and thermal neutrons. A novel event discrimination algorithm, using the properties of the TimePix3 camera, enabled the separation of neutron signatures from the gamma-ray background.
@@ 1 Introduction
Scintillators are a widely used solution for detecting radiation, offering high flexibility and a broad range of applications. The first plastic scintillator, polystyrene doped with m-terphenyl, was discovered by Schorr and Torney in 1950 [1]. Since then, a large variety scintillating materials have been discovered and used in particle physics, medicine, biology and other scientific fields in combination with visible light detectors [2].
@@
sees both a neutron and a positron being produced [6]. In a scintillator with a short decay time, the time difference between the prompt scintillation light from the positron capture and delayed light from the neutron capture gives a clear signature of neutrino detection; this is a role to which plastic scintillators are well suited. Neutron-sensitive plastic scintillators have also found important applications in nuclear security and forensics. Neutrons are a key signature of fissile materials’ decay, and their presence could pose major security threats if used for aggressive purposes; for this reason, there is extensive interest in neutron detection for homeland security and non proliferation (e.g. [7, 8, 9]). Furthermore, the resilience of plastic scintillators to changes in temperature and humidity allows their use in a wide variety of climatic conditions such as ports and countries border crossings [2].
Maximizing the performance of plastic scintillators often demands complex geometries, which are difficult to achieve using traditional manufacturing techniques. In this article, we present a more versatile approach based on additive manufacturing for producing neutron-sensitive plastic scintillators. We also detail strategies to integrate these scintillators into a functional detector system employing a fast optical camera.
@@ 2.2 Additive Manufacturing of Scintillators
Additive manufacturing, or 3D-printing, offers a way to solve the problems outlined in Section 2.1. It allows users to cheaply and rapidly produce scintillators in a scalable manner, and enables easy customization of the composition, internal geometry and external geometry of the scintillator. This is particularly true for plastic scintillators, which can be produced using off-the-shelf equipment. Two main methods have previously been used to 3D-print scintillators: stereolithography (SLA) and fused deposition modelling (FDM) [14, 15]. Both work on the basic principle of building up a piece, layer by layer. In SLA, an ultraviolet (UV) laser is used to selectively cure a liquid resin. This method allows for easy inclusion of additives such as dyes, fluorophores or neutron-sensitive isotopes. It also offers a high printing resolution, with the smallest feature possible being set by the size of the UV laser spot. However, it can only be used with a limited subset of plastics; the resins that can be used also pose a significant health hazard.
@@
FDM, meanwhile, extrudes a thermoplastic filament through a heated nozzle. FDM printing can be used with a wider variety of plastics than SLA, enabling the direct printing of scintillators without additives. It also allows multiple materials or filaments to be printed together. FDM printers are also available at lower cost. The production of customized filaments requires a separate system; as a result, its use for scintillators has been explored less thoroughly than the use of SLA. Early experiments by Hamel and Lebouteiller [19] used poly-lactic acid doped with PPO and 1,4-Bis (2-methylstyryl) benzene (bis-MSB) to print scintillators, to moderate results. More successfully, Berns et al. [20, 21] have demonstrated the ability to print a polystyrene-based scintillator with FDM, while Sibilieva et al. [22] have used FDM to print inorganic scintillators in a plastic matrix. The work of Berns et al. has recently been extended by Li et al. [23], using FDM to produce a segmented detector for particle tracking.
@@ 3 Neutron Detection
Neutrons are difficult to detect directly. Instead, indirect methods have to be used. One option is to use activation foils, made from isotopes that will be transmuted to a radioactive isotope when they absorb a neutron [24]. By measuring the radiation output from these foils post-exposure to a neutron source, the neutron flux from the source can be inferred. This is an effective, and easy-to-implement method, but does not allow for real-time measurements of the source flux. A more real-time method is to measure protons recoiling from neutron impact [25]. This requires the use of a material with a large proportion of hydrogen, making it suitable for use with plastic scintillators. However, the proton signal can be hard to distinguish from gamma rays and other forms of radiation, limiting its utility.
@@
This reaction produces 4.78 MeV of energy, which is divided between the two reaction products, with MeV and MeV [27]. This energy is considerably higher than that from reactions [2]. It also produces an alpha particle rather than a gamma-ray, meaning that the scintillator light will be emitted close to the interaction. For thermal neutrons, the cross-section of this reaction is 940 barns, giving a strong sensitivity to thermal neutrons. Lithium compounds can easily be integrated into a scintillator, making this interaction a highly suitable method for producing a neutron-sensitive plastic scintillator.
@@ 4.1 Filament Production
The polystyrene was combined with PTP as a primary fluorophore and POPOP as a secondary fluorophore, with 2% of the former and 0.05% of the latter being found to give maximum light output. A pure polystyrene filament proved to be too brittle to effectively 3D-print. It would snap when exposed to the rapid movements of the print head, or be crushed in the extruder, leading to print jams. To increase the flexibility of the filament, a plasticizer was added. Biphenyl, which dissolves well in polystyrene, was chosen. The minimum necessary biphenyl content was found to be 7.5%, which gave the flexibility required to print effectively without significant impact on light output. Finally, to allow the scintillator to detect neutrons, 0.1% by weight of was added to the material, in the form of powder. All materials were procured from Sigma Aldrich.
@@
The filament obtained, as described in Section 4.1, was printed using a Prusa i3 3D-printer [30], as seen in Figure 2(b). Most 3D-printed objects use a shell around an infill structure. However, this is unsuitable for printing transparent objects like scintillators, as the infill structure traps gas bubbles. To avoid this, the printing parameters had to be modified to ensure maximum transparency. 100% infill was selected, to ensure that the final scintillator was filled completely with plastic. To reduce surface roughness, especially on the sides, a layer height of 0.05 mm was selected. The infill layers were also directly aligned, to minimise gas bubbles. There is a strong relationship between printing temperature and transparency, with higher temperatures producing a more transparent scintillator. However, higher temperatures also reduce the shape accuracy of the printing, as the plastic can more easily flow before cooling. A print temperature of 255∘ C, with no cooling of the print, was found to give the highest transparency while retaining acceptable shape tolerances. To ensure the extruded plastic had time to fill all space on the layer below, the print speed was set to 12 mms-1, 20% of the basic setting. The flow rate was also increased, to 112% of the base flow rate. This combination of settings was found to give an effectively transparent scintillator. However, there were imperfections in the printing of some scintillators, caused by the presence of gas bubbles, or by underextrusion. This was due to inconsistencies in the diameter of the filament. Using a filament extruder with a fixed 1.75 mm nozzle, would avoid this issue.
@@
Scintillators were printed in a variety of shapes and thicknesses. While all displayed sensitivity to and radiation, low sensitivity to neutrons was observed. Simulations indicated that this was due to low lithium content. However, increasing the lithium content would result in decreased transparency, reducing effective light output for a single-material scintillator. This was achieved by printing a two-part scintillator, with a cavity in the centre. This cavity could be packed with powder, increasing the lithium content of the scintillator, but retaining high transparency.
@@ 5 Detector Performance
The scintillator was exposed to radioactive sources including 90Sr, 137Cs and 252Cf, allowing for tests of the its response to and neutrons. The scintillating block was placed 10 cm from the sources, in line with the camera lens. The 90Sr and 137Cs source were used with no additional items in the light-tight box. The californium source, meanwhile, was contained in a 5 cm thick lead jacket, with a 10 cm High-Density Poly-Ethylene (HDPE) thermalizer between it and the scintillator. When a combination of and neutrons was required, the 137Cs source was also placed behind the thermalizer, alongside the 252Cf source. These configurations were simulated using the Geant4 Monte Carlo package, which showed that the thermalizer and lead jacket would create a neutron spectrum with two components; a fast component with energies at around 1 MeV and a slow component with energies at around 0.03 eV [36]. The fast component dominated, with 70% of the neutrons, but the slow component was still significant, containing 20% of the neutrons, allowing us to test the scintillator’s sensitivity to both the fast and thermal neutron components.
Furthermore, observations of the 90Sr source were used to determine the optimal composition of the scintillator, testing concentrations of PTP and POPOP between 0-4% for the former and 0-0.1% for the latter. As described above, maximum light output with minimum opacity was found to be given with a concentration of 2% PTP and 0.05% POPOP. The 90Sr and 137Cs source were then used to determine the detection performance of the scintillator. It was found that the 3D printed materials had a strong sensitivity to both and . The highest light output was found to be . This light output is about 104 times lower than that of NaI. However, the decay time of the emission is significantly shorter [37].
Measurements of 252Cf were carried out with two different configurations of printed scintillator: a cuboid, and one with a cavity packed with 6Li. The characteristic emission of the latter can be seen in Figure 5(a). These were compared with a commercially produced neutron-sensitive scintillator, Eljen Technology’s EJ-420 [38]. All three scintillators displayed sensitivity to neutrons, with a significant increase in photo-emission over the noise level. The EJ-420 scintillator, however, was found to be more sensitive than the 3D-printed cuboid, with 4.5 times more detections. The scintillator with the increased Li content was, as expected, more sensitive than the cuboid, increasing detections by a factor of 2.
@@ 5.1 Neutron-Gamma Discrimination
The use of the TPX3cam allows a method for neutron-gamma discrimination, without having to tune the scintillator for pulse-shape discrimination. As noted in Section 3, a neutron interaction with will release more energy than that carried by the typical gamma-ray. As such, when the alpha particle and tritium nucleus are detected within the scintillator, a brighter pulse will be emitted. At a basic level, a neutron interaction within the scintillator will cause the emission of a minimum of 143 photons. For gamma rays with a comparable energy to the alpha particle, the photons will be emitted over a broader area, resulting in a more diffuse pulse. The compact, bright pulse from the detection of an alpha particle will be visible as a large cluster in the TimePix output. By identifying the signature of these larger clusters, we can use this to count individual neutron detections. A cluster is defined as a number of hits within a radius of three TimePix pixels within a time resolution of 100 ns; this was chosen as it matches the decay time of the phosphors in the image intensifier. An example of these clusters can be seen in Figure 5(b). By summing the total photon energy for each cluster, we can use this as a proxy for the energy of the incoming particle. In the case of the TimePix, this is in the form of a Time over Threshold (ToT), in nanoseconds.
@@
This was tested with Geant4 simulations of the scintillator, using both thermal neutrons and 662 keV gamma rays. Photons leaving the scintillator and crossing a collector plane were counted, and assigned to clusters, using both time and spatial correlations. The total photon energy was plotted against the total number of collected photons. This showed a clear difference between gamma rays and neutrons, with the neutrons having a ’tail’ containing significantly more events, as can be seen in Figure 6. This result was also obtained with the 3D printed scintillator, using the 137Cs and 252Cf sources to produce gamma-only and neutron-only signals. The results of this can be seen in Figure 7, which shows a clear cut-off in the gamma-ray data at 5 hits per cluster. By setting a minimum value at 5 hits, we are thus able to draw a distinction between neutrons and gamma rays. Based on this, we can plot a cluster energy spectrum for both simulated data and for the real scintillator. Both show a similar peak, associated with thermal neutrons, shown in Figure 8.
This method, as it relies on the neutron-6Li reaction, is not sensitive to fast neutrons. These can be detected by the scintillator, through proton recoil. However, simulations have shown that proton recoil does not cause the same scintillation signature to which this method is sensitive. As such, this method does not allow for the detection of fast neutrons.

### v2

@@ Abstract
This work reports on the performance of a novel neutron-sensitive scintillating detector fabricated using Fused-Deposition Modelling (FDM) additive manufacturing. FDM is a cost-effective 3D-printing method employing flexible plastic filaments to create custom-shaped components. Scintillating filaments, based on polystyrene doped with p-terphenyl and 1,4-bis (5-phenyloxazol-2-yl) benzene, and enriched with 6LiF to enable neutron sensitivity were manufactured in house and achieved visible scintillation with a light output of 30 5 photons per MeV. Printed scintillators (both a 2 cm 2 cm 0.5 cm cuboid and a 2 cm 2 cm 1 cm cuboid containing a 6LiF-packed cavity) were then integrated into a detector system consisting of an image intensified TimePix3 camera, offering high spatial and temporal resolution. The detector performance was compared with Geant4 simulations of the scintillating sensor’s response to electrons, gamma-rays, and thermal neutrons. A novel event discrimination algorithm, using the properties of the TimePix3 camera, enabled the separation of neutron signatures from the gamma-ray background.
@@ 1 Introduction
Scintillators are a widely used solution for detecting radiation, offering high flexibility and a broad range of applications. The first plastic scintillator, polystyrene doped with m-terphenyl, was discovered by Schorr and Torney in 1950 [1]. Since then, a large variety of scintillating materials have been discovered and used in particle physics, medicine, biology and other scientific fields in combination with visible light detectors [2].
@@
sees both a neutron and a positron being produced [6]. In a scintillator with a short decay time, the time difference between the prompt scintillation light from the positron capture and delayed light from the neutron capture gives a clear signature of neutrino detection; this is a role to which plastic scintillators are well suited. As such, they have been used in a number of experiments, such as ISMRAN, VIDARR and the SuperFGD detector for the T2K experiment [7, 8, 9]. Neutron-sensitive plastic scintillators have also found important applications in nuclear security and forensics. Neutrons are a key signature of fissile materials’ decay, and their presence could pose major security threats if used for aggressive purposes; for this reason, there is extensive interest in neutron detection for homeland security and non proliferation (e.g. [10, 11, 12]). Furthermore, the resilience of plastic scintillators to changes in temperature and humidity allows their use in a wide variety of climatic conditions such as ports and countries’ border crossings [2].
Maximizing the performance of plastic scintillators often demands complex geometries, which are difficult to achieve using traditional manufacturing techniques. In this article, we present a more versatile approach based on additive manufacturing for producing neutron-sensitive plastic scintillators. We also detail strategies to integrate these scintillators into a functional detector system employing a fast optical camera. This detector system is widely applicable, but is particularly targeted at nuclear security applications.
@@ 2.2 Additive Manufacturing of Scintillators
Additive manufacturing, or 3D-printing, offers a way to solve the problems outlined in Section 2.1. It allows users to cheaply and rapidly produce scintillators in a scalable manner, and enables easy customization of the composition, internal geometry and external geometry of the scintillator. This is particularly true for plastic scintillators, which can be produced using off-the-shelf equipment. Two main methods have previously been used to 3D-print scintillators: stereolithography (SLA) and fused deposition modelling (FDM) [17, 18]. Both work on the basic principle of building up a piece, layer by layer. In SLA, an ultraviolet (UV) laser is used to selectively cure a liquid resin. This method allows for easy inclusion of additives such as dyes, fluorophores or neutron-sensitive isotopes. It also offers a high printing resolution, with the smallest feature possible being set by the size of the UV laser spot. However, it can only be used with a limited subset of plastics; the resins that can be used also pose a significant health hazard. SLA printing is also limited in its ability to produce large parts, due to the limited print volumes of available SLA printers.
@@
FDM, meanwhile, extrudes a thermoplastic filament through a heated nozzle. FDM printing can be used with a wider variety of plastics than SLA, enabling the direct printing of scintillators without additives. It also allows multiple materials or filaments to be printed together. FDM printers are also available at lower cost. FDM is also much more suitable for printing larger parts, as FDM is an easily scalable technology; consumer FDM printers can reach volumes of greater than 600,000 cm3 [22]. The production of customized filaments requires a separate system; as a result, its use for scintillators has been explored less thoroughly than the use of SLA. Early experiments by Hamel and Lebouteiller [23] used poly-lactic acid doped with PPO and 1,4-Bis (2-methylstyryl) benzene (bis-MSB) to print scintillators, to moderate results. More successfully, Berns et al. [24, 25] have demonstrated the ability to print a polystyrene-based scintillator with FDM, while Sibilieva et al. [26] have used FDM to print inorganic scintillators in a plastic matrix. The work of Berns et al. has recently been extended by Weber et al. [27], using FDM to produce a segmented detector for particle tracking. This detector was characterized in beam tests by Li et al. [28], showing comparable light output to similar detectors produced from cast plastic scintillators.
@@ 3 Neutron Detection
Neutrons are difficult to detect directly. Instead, indirect methods have to be used. One option is to use activation foils, made from isotopes that will be transmuted to a radioactive isotope when they absorb a neutron [29]. By measuring the radiation output from these foils post-exposure to a neutron source, the neutron flux from the source can be inferred. This is an effective, and easy-to-implement method, but does not allow for real-time measurements of the source flux. A more real-time method is to measure the energy deposited by protons recoiling from neutron impact within the material [30]. This requires the use of a material with a large proportion of hydrogen, making it suitable for use with plastic scintillators. However, the proton signal can be hard to distinguish from gamma rays and other forms of radiation, requiring the use of pulse-shape discrimination techniques to detect it. PSD is a highly effective technique, but is not possible with all plastic scintillators, requiring high concentrations of fluorophores [2].
@@
This reaction produces 4.78 MeV of energy, which is divided between the two reaction products, with MeV and MeV [32]. This energy is considerably higher than that from reactions, though quenching within the scintillator will reduce the light output from these significantly [2]. It also produces an alpha particle rather than a gamma-ray, meaning that the scintillator light will be emitted close to the interaction. For thermal neutrons, the cross-section of this reaction is 940 barns, giving a strong sensitivity to thermal neutrons. Lithium compounds can easily be integrated into a scintillator, making this interaction a highly suitable method for producing a neutron-sensitive plastic scintillator.
@@ 4.1 Filament Production
The polystyrene was combined with PTP as a primary fluorophore and POPOP as a secondary fluorophore. The total photons produced by each scintillator were counted using the detector system described in Section 4.3, with maximum light output being found to be 30 5 photons/MeV with with 2% of PTP and 0.05% of POPOP. A pure polystyrene filament proved to be too brittle to effectively 3D-print. It would snap when exposed to the rapid movements of the print head, or be crushed in the extruder, leading to print jams. To increase the flexibility of the filament, a plasticizer was added. Biphenyl, which dissolves well in polystyrene, was chosen. The minimum necessary biphenyl content was found to be 7.5%, which gave the flexibility required to print effectively without significant impact on light output. Finally, to allow the scintillator to detect neutrons, 0.1% by weight of was added to the material, in the form of powder. All materials were procured from Sigma Aldrich [34].
@@
The filament obtained, as described in Section 4.1, was printed using a Prusa i3 3D-printer [36], as seen in Figure 2(b). Most 3D-printed objects use a shell around an infill structure. However, this is unsuitable for printing transparent objects like scintillators, as the infill structure traps gas bubbles. To avoid this, the printing parameters had to be modified to ensure maximum transparency. 100% infill was selected, to ensure that the final scintillator was filled completely with plastic. To reduce surface roughness, especially on the sides, a layer height of 0.05 mm was selected. The infill layers were also directly aligned, to minimise gas bubbles. There is a strong relationship between printing temperature and transparency, with higher temperatures producing a more transparent scintillator. However, higher temperatures also reduce the shape accuracy of the printing, as the plastic can more easily flow before cooling. A print temperature of 255∘ C, with no cooling of the print, was found to give the highest transparency while retaining acceptable shape tolerances. To ensure the extruded plastic had time to fill all space on the layer below, the print speed was set to 12 mms-1, 20% of the basic setting. The flow rate was also increased, to 112% of the base flow rate. This combination of settings was found to give an effectively transparent scintillator.
However, there were imperfections in the printing of some scintillators, caused by the presence of gas bubbles, which in turn resulted from underextrusion of the plastic. This was due to inconsistencies in the diameter of the filament. These gas bubbles were the main defect noticed during the printing process, and their incidence depended heavily on the quality of the filament batch used for printing, occurring in between 10-25% of the printed scintillators. Using a filament extruder with a fixed 1.75 mm nozzle, would avoid this issue. Otherwise, it proved to be a reliable method for producing scintillators.
@@
Scintillators were printed in a variety of shapes and thicknesses. While all displayed sensitivity to and radiation, low sensitivity to neutrons was observed. Simulations indicated that this was due to low lithium content. However, increasing the lithium content would result in decreased transparency, reducing effective light output for a single-material scintillator. This was achieved by printing a two-part scintillator, with a cavity in the centre. This cavity could be packed with powder, increasing the lithium content of the scintillator, but retaining high transparency. The chosen design used a 2 cm by 2 cm by 1 cm cuboid, with a 1 cm by 1 cm by 0.1 cm cavity 0.5 cm inside the scintillator.
@@ 5 Detector Performance
The scintillator was exposed to radioactive sources including 90Sr, 137Cs and 252Cf, allowing for tests of the its response to and neutrons. The scintillating block was placed 10 cm from the sources, in line with the camera lens. The 90Sr and 137Cs source were used with no additional items in the light-tight box. The 137Cs source was used to measure the light yield of the scintillator, counting total photon emission from the scintillator over a fixed exposure time. The 252Cf source, meanwhile, was contained in a 5 cm thick lead jacket, with a 10 cm High-Density Poly-Ethylene (HDPE) thermalizer between it and the scintillator. A 10 cm lead block was also used to reduce the numbers of gamma rays reaching the scintillator. When a combination of and neutrons was required, the 137Cs source was also placed behind the thermalizer, alongside the 252Cf source. These configurations were simulated using the Geant4 Monte Carlo package, which showed that the thermalizer and lead jacket would create a neutron spectrum with two components; a fast component with energies at around 1 MeV and a slow component with energies at around 0.03 eV [42]. The fast component dominated, with 70% of the neutrons, but the slow component was still significant, containing 20% of the neutrons, allowing us to test the scintillator’s sensitivity to both the fast and thermal neutron components. These simulations, along with all others carried out for this project, used Geant4 11.2, using the FTFP_BERT_HP physics list, with the G4RadioactiveDecayPhysics and G4OpticalPhysics options.
Furthermore, observations of the 90Sr source were used to determine the optimal composition of the scintillator, testing concentrations of PTP and POPOP between 0-4% for the former and 0-0.1% for the latter. As described above, the maximum light output with minimum opacity was found to be given with a concentration of 2% PTP and 0.05% POPOP. The 90Sr and 137Cs source were then used to determine the detection performance of the scintillator. It was found that the 3D printed materials had a strong sensitivity to both and . The highest light output was found to be . This light output is about 104 times lower than that of NaI. However, the decay time of the emission is significantly shorter [43]. The light output achieved is comparable to that achieved with undoped 3D-printed scintillators by Berns et al. [25], but is significantly worse than commercially available scintillators produced using traditional methods; for example, Eljen’s EJ-270 reaches 4,800 [44].
Measurements of 252Cf were carried out with two different configurations of the printed scintillator: a 2 cm 2 cm 0.5 cm cuboid (as seen in Figure 3, and one with a cavity packed with 6Li using the design described in Section 4.2. The characteristic emission of the latter can be seen in Figure 5(a). These were compared with a commercially produced neutron-sensitive scintillator, Eljen Technology’s EJ-420 [45]. All three scintillators displayed sensitivity to neutrons, with a significant increase in photo-emission over the noise level. The EJ-420 scintillator, however, was found to be more sensitive than the 3D-printed cuboid, with 4.5 times more detections. The scintillator with the increased Li content was, as expected, more sensitive than the cuboid, increasing detections by a factor of 2.
@@ 5.1 Neutron-Gamma Discrimination
The use of the TPX3cam allows a method for neutron-gamma discrimination, without having to tune the scintillator for pulse-shape discrimination. As noted in Section 3, a neutron interaction with will release more energy than that carried by the typical gamma-ray. As such, when the alpha particle and tritium nucleus are detected within the scintillator, a brighter pulse will be emitted. Given the total reaction energy of 4.78 MeV [32] and the scintillator response of 30 photons per MeV, 140 photons will be produced for each neutron reaction; though this naive analysis neglects quenching, which will reduce this number significantly. Even so, there will be a bright, compact pulse emitted in the area where the neutron is detected. For gamma rays with a comparable energy to the reaction products, the photons will be emitted over a broader area, resulting in a more diffuse pulse. The compact, bright pulse from the detection of a neutron will be visible as a large cluster in the TimePix output. By identifying the signature of these larger clusters, we can use this to count individual neutron detections. A cluster is defined as a number of hits within a radius of three TimePix pixels within a time resolution of 100 ns; this was chosen as it matches the decay time of the phosphors in the image intensifier. An example of these clusters can be seen in Figure 5(b). By summing the total photon energy for each cluster, we can use this as a proxy for the energy of the incoming particle. In the case of the TimePix, this is in the form of a Time over Threshold (ToT), in nanoseconds.
@@
This was tested with Geant4 simulations of the scintillator, using both thermal neutrons and 662 keV gamma rays. Photons leaving the scintillator and crossing a collector plane were counted, and assigned to clusters, using both time and spatial correlations. The total photon energy was plotted against the total number of collected photons. This showed a clear difference between gamma rays and neutrons, with the neutrons having a ’tail’ containing significantly more events, as can be seen in Figure 6. This result was also obtained with the 3D printed scintillator, using the 137Cs and 252Cf sources to produce gamma-only and neutron/gamma signals, using a 10 cm thick lead block to reduce the number of 252Cf gamma rays reaching the scintillator. The results of this can be seen in Figure 7, which shows a clear cut-off in the gamma-ray data at 5 hits per cluster. By setting a minimum value at 5 hits, we are thus able to draw a distinction between neutrons and gamma rays. Based on this, we can plot a cluster energy spectrum for both simulated data and for the real scintillator. Both show a similar peak, associated with thermal neutrons, shown in Figure 8. 252Cf fission produces multiple neutrons, which poses only a slight challenge to this method of discrimination. Multiple neutrons arriving simultaneously in widely spaced areas of the scintillator will be detected as separate clusters and counted separately. However, if multiple neutrons arrive in the same location, they will be detected as a single cluster and counted together. This will reduce the number of neutrons counted compared to those detected by the scintillator.
This method, as it relies on the neutron-6Li reaction, is not sensitive to fast neutrons, as this reaction has only a small cross-section to fast neutrons. Instead, these must be detected using by measuring proton recoil within the scintillator. However, the signature of proton recoil cannot easily be discriminated from gamma ray interactions within the scintillator without changes in composition to ease pulse-shape discrimination using traditional methods.